# EarthMarker: Visual Prompt Learning for Region-level and Point-level Remote Sensing Imagery Comprehension

### v2

@@ III-A Model Architecture
One challenge in the RS domain is the absence of a visual prompts learning framework to endow MLLMs with fine-grained image understanding capabilities, blocking more complex reasoning. To address this challenge, EarthMarker is proposed, utilizing visual prompting for multi-granularity RS imagery comprehension. As illustrated in Fig. 2, EarthMarker contains four core components: a sharing visual encoding mechanism, a modality-align projection layer, a text tokenizer module, and a LLM decoder. These components work together to deal with multi-modal information, such as text instruction, images, and diverse visual prompting marks including bounding boxes and points, allowing LLM to generate accurate text responses. Each part is introduced as follows in detail.
@@
In particular, the images and corresponding visual prompts share a visual encoding mechanism for feature sharing, enabling the visual encoders to better understand and associate the relationship between images and visual prompts. Specifically, the Mixture of Visual Experts (MoV) [35] is designed to encode the visual information. The MoV incorporates two visual encoders, DINOv2-ViT L/14 [36] and CLIP-ConvNeXt [37], which are pre-trained on distinct network architectures (ViT and CNN), thus offering complementary visual semantics. To refine the robust multi-scale visual features, the input images are downsampled to different resolutions denoted as and then respectively fed into the MoV module to encode. Leveraging the strengths of various visual backbones, visual perception is enhanced and key details in images are refined. Subsequently, the encoded visual features are transformed to the same dimension and concatenated channel-wisely to obtain the integrated multi-scale feature maps represented as . This process can be formulated as
@@
Notably, a key step to the encoder-sharing mechanism is the “Visual Prompt as Images”. Especially, the dimension () of the visual prompts is processed to the same dimension () with the images. Then, the transformed visual prompts P also can be fed into MoV together with the images, the encoded visual prompts expressed as . Similarly, this process is written as
@@
Subsequently, the modality alignment projection layer transforms the visual tokens into the language semantic space. Meanwhile, the text instructions are processed by the tokenizer module, which handles text tokenization and embedding, converting them into text embeddings . After obtaining the projected image tokens, visual prompts tokes, and text instructions embeddings, they are integrated into an entire multi-modal input sequence. LLM decoder takes the multi-modal inputs and generates the response sequence Y, which can be formulated as
It should be noted that we employ Llama 2, a transformer-based decoder-only LLM, as the LLM decoder.
III-B Cross-domain Phased Training
@@
To realize the fundamental image-level understanding, spatial perception, and region/point-level RS data interpretation ability, the cross-domain phased training method is designed. The entire training process is divided into three phases including multi-domain image-text alignment, spatial perception tuning, and RS visual prompting tuning stage. Throughout the training, we keep lightweight training and avoid expensive full-parameters tuning. Furthermore, the disjoint parameters strategy is proposed, namely, the updated parameters of each stage are different. This strategy is conducive to the step-by-step solid understanding of images, and naturally solves the interference between image-text understanding, visual prompting comprehension, and fine-grained instruction-following.
@@
Multi-domain Image-text Alignment. The first phase employs a multi-domain image-text alignment strategy. In this stage, both natural and RS domain image-level data are leveraged for pre-training to bring visual and text knowledge into alignment within a high-dimensional feature space. This strategy enables EarthMarker to deeply understand the holistic semantics of images. Specifically, we utilize the natural scene caption dataset COCO Caption [38], alongside the RS image caption and scene classification subset from the newly constructed RSVP. During this training phase, multi-scale visual features and language representations are integrated into the LLM to develop image-level comprehension capabilities. The MoV module is kept frozen throughout the training, so as to concentrate on refining robust visual features. Only the alignment projection layer, which acts as the visual-language connector, undergoes parameter updates to enhance the multimodal capabilities of the proposed EarthMarker and ensure seamless integration of visual and textual information.
@@
Spatial Perception Tuning. In the previous step, EarthMarker achieved image-level comprehension capability. In this step, to acquire spatial perception and object-level comprehension, the nature scene publicly available datasets, RefCOCO [39] and RefCOCO+ [40] are transformed into the instruction-following format. Throughout the training, the attention layers of LLM are unfrozen for aligning spatial region features with language embeddings. Specifically, LLM’s key module self-attention head is composed of key , query , and the value , which are transformed by several linear layers. The -th implementation equations can be expressed as follows
@@
RS Visual Prompting Tuning. The last stage focuses on accurately following user instructions and achieving complex region-level and point-level visual reasoning tasks. The MoV, alignment projection, and LLM are fixed. The LoRA method is adopted for tuning. We load the weights trained in the previous phase and continue training EarthMarker on RSVP-3M region-text and point-text parings, which contain the fine-grained referring object classification and referring brief caption data. Specifically, several learnable low-rank adapter matrices , , and are inserted into Transformer layers of LLM. The is the number of attention heads. The adapted multi-head attention is denoted as , thus the output of the -th adapted Transformer attention is formulated as
In conclusion, the cross-domain phased training endows EarthMarker with various granular (e.g., image-level, point-level, and region-level) multimodal instruction capabilities in the RS domain. In the first multi-domain image-text alignment stage, the LLM is efficiently converted into an MLLM, which is capable of image-level understanding. Subsequently, by utilizing the nature scene referring datasets, EarthMarker is equipped with the fundamental spatial perception of images. This is beneficial for subsequent developments of referring ability in the RS domain. Furthermore, by leveraging the RS visual prompting datasets RSVP-3M, EarthMarker is endowed with image understanding at both the region level and point level. Notably, different field’s datasets are adopted for training, and enhancing open-vocabulary reasoning ability. It should be emphasized that during the whole training, our updatable parameters are disjoint, preventing interference between understanding images at different granularity and the capability to follow visual prompts.
@@ IV RS Visual Prompting Dataset Construction
In this section, a visual prompting dataset named RSVP-3M is presented. RSVP-3M is the first visual prompting instruction dataset in the RS field, designed to advance image-level and fine-grained point-level, region-level RS MLLMs. Specifically, RSVP-3M contains over 3 million multimodal dialogue data with visual prompting marks. Those multi-granularity visual prompting data are restructured and cleaned from existing publicly available RS datasets. Furthermore, the GPT-4V [18] is employed for automatic annotation to construct a high-quality complex visual reasoning dataset[15]. A detailed explanation of the construction of the RSVP-3M dataset is introduced as follows.
@@ IV-A Data Conversion and Annotation from Public RS Datasets
A part data of the dataset RSVP-3M is constructed by restructuring and relabeling existing RS datasets. A range of visual task types is covered, containing image classification, instance segmentation, object detection, image caption, and region caption, see Tab. I. The image-level, region-level, and point-level data are derived from different RS datasets. Firstly, image-level visual prompting data is converted from image classification and captioning datasets. For the two type datasets, image-level visual instructions are used, with the bounding box serving as the visual prompt to obtain the image’s category or brief caption. Subsequently, the region-level data is based on object detection datasets. The ground truth bounding boxes are used as visual prompts to guide the model to identify the object-level or region-level categories accurately. Additionally, the point-level data is transformed from segmentation datasets. For instance segmentation, the representative points extracted from masks corresponding to instances are used as point-level visual prompts. For semantic segmentation, each image is divided into patches, and the points are randomly sampled within each patch as the visual prompts, with the category retrieved from the corresponding segmentation map.
Tasks | Raw Data | Samples |
Image Captioning | NWPU-Captions [41] | 169,981 |
RSITMD[42] | 24,387 | |
Sydney-Captions [43] | 2,837 | |
Region Captioning | DIOR-RSVG [19] | 31,491 |
Scene Classification | NWPU-RESISC45 [44] | 94,500 |
OPTIMAL 31 [45] | 3,720 | |
RSD46 [46] | 350,685 | |
WHU-RS19 [47] | 3,015 | |
Referring Object Classification (Box) | DOTA V2 [48] | 99,774 |
FAR1M [49] | 17,074 | |
NWPUVHR10 [50] | 1,888 | |
RSOD [46] | 1,465 | |
UCAS-AOD [51] | 3,203 | |
VisDrone [52] | 128,531 | |
MAR20 [53] | 2,096 | |
DOSR [54] | 1,127 | |
LEVIR [55] | 4,633 | |
HRSC2016 [56] | 2,171 | |
HRRSD [57] | 12,519 | |
Referring Object Classification (Point) | Vaihingen [58] | 45,104 |
Potsdam [58] | 504,139 | |
Hi-UCD [59] | 125,908 | |
NWPUVHR10 [50] | 956 | |
isAID [60] | 294,355 | |
UAVid [61] | 1,167,543 | |
SOTA [62] | 171,583 | |
FAST [62] | 263,056 | |
WHU [63] | 33,022 | |
Relationship Analyses | HRRSD [57] | 11,602 |
RSOD [46] | 934 | |
NWPUVHR10 [50] | 1,950 | |
LEVIR [55] | 3,956 | |
UUCAS-AOD [51] | 2,892 | |
DIOR-RSVG [19] | 15,274 | |
Grounded Captioning | DIOR-RSVG [19] | 15,274 |
GeoChat [4] | 35,264 | |
NWPUVHR10 [50] | 975 |
@@
In RSVP-3M, each data item consists of visual prompts, user instructions, and image. The visual prompts in the user instructions or model answers are expressed as or . For the point-level data, for example, the user instruction which guides referring object classification is “Please identify the category of each marked point in the image”. The answer format is “‘points’”. In addition, for the region-level data, take the scene of the airport as an example. The user instruction for airport region captioning is “Please provide the brief caption of each marked region in the image, and the corresponding answer format generated by the model is “ ‘bbox’”. The data structures of other visual tasks are similar to those explained above. Through the transformation and re-annotation based on public datasets, the visual prompting dataset RSVP-3M is effectively developed, featuring image-point-text and image-region-text pairings.
@@ IV-B GPT4V-assisted Visual Prompting Data Generation
The aforementioned public datasets only provide simple classification information and brief captions, which are insufficient for intelligent interpretation of complex RS imagery. To mitigate the limitation and develop a more detailed and explicit RS visual prompting dataset, the language prompts for GPT-4V are carefully crafted for generating data featuring various complex visual reasoning. The complex fine-grained visual tasks involve detailed image captioning, inter-relationship analysis, and grounding captioning. We adopt the Set-of-Marks[64] (SoM) prompting, which can effectively unleash the extraordinary visual grounding ability of GPT-4V, to obtain comprehensive and unique characteristics from the RS imagery.
The data generated using GPT-4V not only compensates for the lack of information in brief captions but also provides detailed descriptions that reveal the spatial and semantic relationships between different regions in the image. For example, in aerial imagery, it is feasible to identify the general category of the image and provide a simple description. Additionally, detailed descriptions, such as the spatial layout of tennis courts, basketball courts, playgrounds, the relationships among these areas, and the activities of people on the playground, can be conducted. The RSVP-3M dataset, supplemented by public datasets and data generated by GPT-4V, covers a wide range of fine-grained visual reasoning tasks, enhancing the richness and diversity of the data.
@@ V Experiments
In this section, we present extensive experiments to validate the superior performance of EarthMarker. In Section V-A, we introduce the implementation details. Subsequently, we conduct qualitative and quantitative analyses to provide a holistic view of EarthMarker’s performance from Sections V-B to V-E.
V-A Implementation Details

### v3

@@ III-A Model Architecture
In order to set up a flexible and effective interpretation way on complicated RS data, visual prompting-based MLLM is proposed. As illustrated in Fig. 2 (a), EarthMarker contains four core components: a sharing visual encoding module, a modality-align projection layer, a text tokenizer module, and an LLM decoder. These components work together to deal with multi-modal information, such as text instruction, images, and diverse visual prompts including bounding boxes and points, allowing LLM to generate accurate text responses. Each part is introduced as follows in detail.
Theoretically, the first key point is establishing unified representation[50, 51, 52] across different modalities (e.g., text, images, and visual prompts), interpretation levels (e.g., image-level, region-level, and point-level), and a wide range of visual tasks. Specifically, all inputs , where , are embedded into the latent space, which is expressed as , where are the respective embedding functions that map each input to a latent representation . Subsequently, the mathematics formula of the unified feature space processing can be expressed as
where is a function that integrates text, image, region, and point-level information into a unified feature representation .
@@
In EarthMarker, different modalities are processed through distinct encoder designs, ultimately achieving unified feature representation. In particular, the images and corresponding visual prompts share a visual encoding mechanism for feature sharing, enabling the visual encoders to better understand and associate the relationship between images and visual prompts. Specifically, the Mixture of Visual Experts (MoV) [53] is designed to encode the visual information. The MoV incorporates two visual encoders, DINOv2-ViT L/14 [54] and CLIP-ConvNeXt [51], which are pre-trained on distinct network architectures, thus offering complementary visual semantics. To refine the robust multi-scale visual features, the input images are downsampled to different resolutions denoted as and then respectively fed into the MoV module to encode. Leveraging the strengths of various visual backbones, visual perception is enhanced and key details in images are refined. Subsequently, the encoded visual features are transformed to the same dimension and concatenated channel-wisely to obtain the integrated multi-scale feature maps represented as . This process can be formulated as
This approach refines the multi-scale visual information, enhancing the model’s perception across different resolutions.
@@
Notably, a key step to the encoder-sharing mechanism is called “Visual Prompts as Images”. Especially, the dimension () of the visual prompts is processed to the same dimension () with the images. Then, the transformed visual prompts P also can be fed into MoV together with the images, the encoded visual prompts expressed as . Similarly, this process is written as
where , , and represents image-level, region-level, and point-level visual prompts, respectively.
@@
Subsequently, the modality alignment projection layer transforms the visual tokens into the language semantic space. Meanwhile, the text instructions are processed by the tokenizer module, which handles text tokenization and embedding, converting them into text embeddings . After obtaining the projected image tokens, visual prompts tokes, and text instructions embeddings, they are integrated into an entire multi-modal input sequence.
To achieve cross-modal multi-granularity knowledge mutual enhancement [15], the LLM decoder is designed as a bridge, linking various knowledge to ensure the model integrates a combination of global, regional, and detailed insights. LLM decoder takes the multi-modal inputs and dynamically adjusts the response, which features flexible task adaptation[53, 55]. Each task is triggered by text instructions , which guide the model in adapting to different interpretation levels. The response process of the developed model is expressed through the following equations.
where indicate whether image-level, region-level, or point-level granularity is used for the task, represent the different levels interpretation results. signifies the proposed model, denotes the model output which dynamically adapts diverse tasks at different interpretation levels.
III-B Cross-domain Training
@@
To bridge the domain gap between natural and RS data, and effectively adapt general-domain knowledge into the RS domain, the cross-domain cross-domain learning training method is designed. As depicted in Fig. 2 (b), the entire training process is divided into three phases including multi-domain image-text alignment, spatial perception tuning, and RS visual prompting tuning stage. Throughout the training, we keep lightweight training and avoid expensive full-parameters tuning. Furthermore, the disjoint parameters strategy is proposed, namely, the updated parameters of each stage are different. This strategy is conducive to the step-by-step solid understanding of images, and naturally solves the interference between image-text understanding, visual prompting comprehension, and multi-modal instruction-following.
@@
Multi-domain Image-text Alignment. The first phase employs a multi-domain image-text alignment strategy. In this stage, both natural and RS domain image-level data are leveraged for pre-training to bring visual and text knowledge into alignment within a high-dimensional feature space. This strategy enables EarthMarker to deeply understand the holistic semantics of images. Specifically, we utilize the natural scene caption dataset COCO Caption [56], alongside the RS image caption and scene classification subset from the newly constructed RSVP. During this training phase, multi-scale visual features and language representations are integrated into the LLM to develop image-level comprehension capabilities. The MoV module is kept frozen throughout the training to concentrate on refining robust visual features. Only the alignment projection layer, which acts as the visual-language connector, undergoes parameter updates to enhance the multimodal capabilities of the proposed EarthMarker and ensure seamless integration of visual and textual information.
@@
Spatial Perception Tuning. In the previous step, EarthMarker achieves image-level comprehension capability. In this step, to acquire spatial perception and object-level comprehension, the nature scene publicly available datasets, RefCOCO [57] and RefCOCO+ [58] are utilized. Throughout the training, the attention layers of LLM are unfrozen for aligning spatial regional features with language embeddings. Specifically, LLM’s key module self-attention head is composed of key , query , and the value , which are transformed by several linear layers. The -th implementation equations can be expressed as follows
@@
RS Visual Prompting Tuning. The last stage focuses on accurately following user visual instructions and achieving complex region-level and point-level visual reasoning tasks. The MoV, alignment projection, and LLM are fixed. The LoRA method is adopted for tuning. We load the weights trained in the previous phase and continue training EarthMarker on RSVP region-text and point-text parings, which contain the fine-grained referring object classification and referring brief caption data. Specifically, several learnable low-rank adapter matrices , , and are inserted into Transformer layers of LLM. The is the number of attention heads. The adapted multi-head attention is denoted as , thus the output of the -th adapted Transformer attention is formulated as
In conclusion, the cross-domain training endows EarthMarker with various granular (e.g., image-level, point-level, and region-level) multi-modal instruction capabilities in the RS domain. In the first multi-domain image-text alignment stage, the LLM is efficiently converted into an MLLM, which is capable of image-level understanding. Subsequently, by utilizing the nature scene referring datasets, EarthMarker is equipped with the fundamental spatial perception of images. This is beneficial for subsequent developments of referring comprehension ability in the RS domain. Most importantly, the training on large-scale natural scenes lays the foundation for EarthMarker to achieve the adaptation to diverse resolutions imagery, and real-world scenarios. Furthermore, by leveraging the RS visual prompting dataset RSVP, EarthMarker is endowed with image understanding across the image, region, and point levels. Notably, different field datasets are adopted for training, and enhancing open-vocabulary reasoning ability. It should be emphasized that during the whole training, our updatable parameters are disjoint, preventing interference between understanding images at different granularity and the capability to follow visual prompts. After step-wise progressive learning, EarthMarker is capable of flexibly shifting focus between global context and fine details as needed.
@@ IV RS Visual Prompting Dataset Construction
Tasks | Raw Datas | Samples | Resolutions |
Image Captioning | NWPU-Captions [59] | 169,981 | 0.230m |
RSITMD[60] | 24,387 | ||
Sydney-Captions [61] | 2,837 | 0.5m | |
Region Captioning | DIOR-RSVG [23] | 31,491 | 0.530m |
Scene Classification | NWPU-RESISC45 [62] | 94,500 | 0.230m |
OPTIMAL 31 [63] | 3,720 | ||
RSD46 [64] | 350,685 | 0.52m | |
WHU-RS19 [65] | 3,015 | m | |
Referring Object Classification (Box) | DOTA V2 [66] | 99,774 | |
FAR1M [67] | 17,074 | 0.30.8m | |
NWPUVHR10 [68] | 1,888 | 0.52m | |
RSOD [64] | 1,465 | 0.33m | |
UCAS-AOD [69] | 3,203 | ||
VisDrone [70] | 128,531 | ||
MAR20 [71] | 2,096 | ||
DOSR [72] | 1,127 | 0.52.5m | |
LEVIR [73] | 4,633 | 0.21m | |
HRSC2016 [74] | 2,171 | 0.42m | |
HRRSD [75] | 12,519 | 0.151.2m | |
Referring Object Classification (Point) | Vaihingen [76] | 45,104 | 0.09m |
Potsdam [76] | 504,139 | 0.05m | |
Hi-UCD [77] | 125,908 | 0.1m | |
NWPUVHR10 [68] | 956 | 0.52m | |
isAID [78] | 294,355 | ||
UAVid [79] | 1,167,543 | ||
SOTA [80] | 171,583 | ||
FAST [80] | 263,056 | 0.30.8m | |
WHU [81] | 33,022 | 0.3m | |
Relationship Analyses | HRRSD [75] | 11602 | 0.151.2m |
RSOD [64] | 934 | 0.33m | |
NWPUVHR10 [68] | 1,950 | 0.52m | |
LEVIR [73] | 3,956 | 0.21m | |
UCAS-AOD [69] | 2,892 | ||
DIOR-RSVG [23] | 15,274 | 0.530m | |
Grounded Captioning | DIOR-RSVG [23] | 15,274 | 0.530m |
GeoChat [6] | 35,264 | ||
NWPUVHR10 [68] | 975 | 0.52m |
In this section, a visual prompting dataset called RSVP is presented. RSVP is the first visual prompting instruction dataset in the RS field. EarthMarker is designed to advance coarse-grained image-level and fine-grained point/region-level RS-specific MLLMs. Specifically, RSVP contains roughly 3.65 million multi-modal dialogue data with visual prompts. Those multi-granularity visual prompting data are restructured and cleaned from existing publicly available RS datasets, covering global data distribution and different spatial resolutions. Furthermore, the GPT4V[22] is employed for automatic annotation to construct a high-quality complex visual reasoning dataset[16]. A detailed explanation of the construction of the RSVP dataset is introduced as follows.
@@ IV-A Data Conversion and Annotation from Public RS Datasets
A part of the data of the dataset RSVP is constructed by restructuring and relabeling existing RS datasets. As shown in Tab. I, a range of visual task types are covered, containing scene classification, referring object classification, image captioning, region captioning, relationship analyses, etc. Notably, RSVP includes various RS images that feature different spatial resolutions, containing very high-resolution (VHR) optical images and medium- to low-resolution imagery. In addition, most of the raw data comes from Google Earth which is sourced from various satellites or aerial photographs, providing broad geographic coverage.
In RSVP, each data item consists of visual prompts, user instructions, and an image. The visual prompts in the user instructions or model answers are expressed as or . Notably, when performing image-level tasks, the covers the entire image. The image-level, region-level, and point-level data are derived from different RS dataset types. For example, image-level visual prompting data is converted from image classification and captioning datasets. For the two type datasets, the bounding box served as the image-level visual prompt, which represents the entire image. For image-level detailed captioning tasks, the text instruction is “Please provide a detailed description of the in the image” to guide the model for image-level captioning.
@@
The region-level data is based on object detection datasets. The ground truth bounding boxes are used as visual prompts to instruct the model to identify the object-level or region-level categories accurately. Take the scene of the airport as an example, the text instruction for the region captioning task is “Please provide the brief caption of each marked region in the image, and the corresponding answer format generated by the model is “ ‘bbox’”.
Notably, the point-level data is transformed from segmentation datasets. For instance segmentation, the representative points extracted from masks corresponding to instances are used as point-level visual prompts. For semantic segmentation, each image is divided into patches, and the points are randomly sampled within each patch as the visual prompts, with the category retrieved from the corresponding segmentation map. The user instruction which guides the model to accomplish point-level referring object classification task is “Please identify the category of each marked point in the image”. The answer format of point-level is “‘points’”.
The data structures of other visual tasks are similar to those explained above. Through the transformation and re-annotation based on public datasets, the visual prompting dataset RSVP is effectively developed, containing image-point-text and image-region-text pairings.
@@ IV-B GPT4V-assisted Visual Prompting Data Generation
The aforementioned public datasets only provide simple classification information and brief captions, which are insufficient for intelligent interpretation of complex RS imagery. To mitigate the limitation and develop a more detailed and explicit RS visual prompting dataset, the language prompts for GPT4V are carefully crafted for generating data featuring various complex visual reasoning. The complex fine-grained visual tasks involve detailed region captioning, object inter-relationship analysis, grounding captioning, etc. We adopt the Set-of-Marks[82] (SoM) prompting, which can effectively unleash the extraordinary visual grounding ability of GPT4V, to obtain comprehensive and unique characteristics from the RS imagery.
The data generated using GPT4V not only compensates for the lack of information in brief captions but also provides detailed descriptions that reveal the spatial and semantic relationships between different regions in the image. For example, in aerial imagery, it is feasible to identify the general category of the image and provide a simple description. Additionally, detailed descriptions, such as the spatial layout of tennis courts, basketball courts, and playgrounds, the relationships among these areas, and the activities of people on the playground, can be conducted. The RSVP dataset, supplemented by public datasets and data generated by GPT4V, covers a wide range of fine-grained visual reasoning tasks, enhancing the richness and diversity of the data.
@@ V Experiments
In this section, we present extensive experiments to validate the superior performance of EarthMarker. In Section V-A, we introduce the implementation setups including evaluation metrics and implementation details. Subsequently, we conduct qualitative analyses to provide a holistic view of EarthMarker’s performance from Sections V-B to V-E.
V-A Implementation Setups
V-A1 Evaluation Metrics
For scene classification tasks, we utilize the scene classification accuracy metric, which specifically measures the proportion of correctly classified scenes out of the total number of scenes. Subsequently, for the evaluation of image and region captioning tasks, we employ a range of metrics designed to assess the quality of language generation. These metrics include BLEU scores[83], which is a modified precision metric with a sentence-brevity penalty, calculated as a weighted geometric mean over different length n-grams. METEOR [84] uses exact, stem, synonym, and paraphrase matches between n-grams to align sentences, before computing a weighted F-score with an alignment fragmentation penalty; ROUGE-L [85] focuses on the longest common subsequence to assess the fluency and relevance of the captions; CIDEr[86], applies term frequency-inverse document frequency (TFIDF) weights to n-grams in the candidate and reference sentences, which are then compared by summing their cosine similarity across n-grams; and SPICE[87], which assesses the quality of image captions, by comparing the generated captions with reference captions the based on semantic similarity. Furthermore, for referring object classification tasks evaluation, following[14, 16], we use Semantic Intersection over Union (S-IOU) and Semantic Similarity (SS)[88]. SS measures the similarity of predicted and ground truth labels in a semantic space, while S-IOU reflects the overlap of words.
V-A2 Implementation Details